# ImmerseDiffusion: A Generative Spatial Audio Latent Diffusion Model

### v1

@@ 1 Introduction
This paper introduces ImmerseDiffusion, a text-conditioned generative spatial audio model designed to produce 3D Ambisonic soundscapes with localization in horizontal, vertical, and distance dimensions, while also accounting for temporal and environmental factors like room size and reverberation. Our model generates FOA-domain audio, and features two operational modes: descriptive and parametric. The descriptive model generates spatial audio based on textual descriptions of sound sources and their spatial and environmental details, suited for narrative-driven applications such as cinematic audio. The parametric model combines textual descriptions of sound sources with numerical spatial parameters, ideal for machine-centered uses like game engines and virtual simulations. To the best of our knowledge, this is the first spatial audio generative model. Consequently, we introduce new metrics to evaluate the quality of the generated FOA audio. These metrics include Ambisonics Fréchet Audio Distance (FAD), spatial Kullback–Leibler (KL) divergence, and spatial CLAP scores. Additionally, we assess spatial accuracy using azimuth, elevation, and distance L1 scores, besides spatial angle which are all based on sound intensity vectors [10].
@@ 2.2 Spatial Codec
To address compression, various approaches have been explored in the literature. Some models leverage autoencoders with quantized latents such as Vector Quantized Variational Autoencoders VQ-VAE (e.g., [12]) or Residual Vector Quantization RVQ-VAE (e.g., [13]). Other models leverage continuous VAE latents to avoid quantization information loss [3, 9, 2]. Following [3], we utilize the 1D convolutional U-Net autoencoder from the Descript Audio Codec (DAC) [14] and replace the discrete RVQ bottleneck with a continuous VAE bottleneck. We also eliminate the decoder’s $tanh()$ activation to prevent harmonic distortion, as noted in [9]. Additionally, we modify the training loss functions to suit the FOA components. Unlike stereo codecs, which leverage losses like Mid-Side Short Time Fourier Transform (MS-STFT) for proper stereo reconstruction, FOA does not need such losses due to the inherent orthogonality of the $X$, $Y$, and $Z$ channels. The training loss function of our FOA codec is as follows:
{dmath}
L_codec = λmrstft4 L_mrstft_W + λmrstft4 L_mrstft_X + λmrstft4 L_mrstft_Y + λmrstft4 L_mrstft_Z + λ_kl L_kl + λ_adv L_adv + λ_fm L_fm
In this context, $\mathcal{L}_{\text{mrstft}}$ represents the Multi-Resolution STFT loss for each channel, collectively contributing to the generation loss. $\mathcal{L}_{\text{kl}}$ denotes KL divergence loss applied to the VAE bottleneck, $\mathcal{L}_{\text{adv}}$ signifies the adversarial (discrimination) loss, while $\mathcal{L}_{\text{fm}}$ refers to the feature matching loss. The $\lambda$ parameters are weights perceptually assigned to adjust each corresponding loss contribution in total loss [3]. Our ambisonic codec model traonsforms a 4-channel FOA signal with length $L$ into $64$ channel latent with length $L/2048$ with an overall compression factor of $128$.
Table 1: Specifications and Evaluation Results of the trained FOA Codecs with continuous VAE latents on Spatial AudioCaps.
Model
@@
1D-Conv U-Net(32X) 1024 128 43Hz 1.69 2.83
1D-Conv U-Net (64X) 1024 64 43Hz 1.70 2.62
1D-Conv U-Net (128X) 2048 64 21.5Hz 1.73 2.55
1D-Conv U-Net-$\beta$ (128X) 2048 64 21.5Hz 2.33 4.13
@@ 2.3 Spatial, Temporal and Environmental Conditioning
ImmerseDiffusion offers spatial control in audio generation via two models: descriptive and parametric. The descriptive model utilizes text-based prompts to define the sound source, its spatial attributes, and environmental acoustics. In contrast, the parametric model uses non-spatial text prompts only for sound source descriptions while providing numerical values for spatial and environmental parameters. Moreover, in line with previous stereo diffusion models  [3, 9], both models employ temporal conditioning, including start time and total time as well. Figure 1 (I) and (II) illustrate how descriptive and parametric conditioning incorporate location and environmental parameters through text or a combination of text and numerical values, along with temporal conditioning.
@@
For the parametric model, we replace ELSA text encoder with the LAION CLAP [5] text encoder, which is solely trained on non-spatial prompts to validate parametric conditioning and avoid potential spatial bias from ELSA embeddings. Alongside text embeddings, this model includes three numerical spatial conditioning parameters (azimuth, elevation, distance) and two environmental acoustic conditioning parameters (room size, 30dB decay time). It also comprises two temporal conditioning parameters (start time, total time), similar to the descriptive model.
@@ 2.4 Diffusion Model
Several generative audio latent diffusion models utilize 1D convolutional U-Net structures for their diffusion blocks (e.g., [2, 3]). However, the Diffusion Transformer (DiT), successful in image generation [22], has recently been adapted for music generation [9, 23]. In our work, we employ a DiT that features a cascade of blocks that include self-attention, cross-attention, and gated MLP components with layer normalization and skip connections, similar to that used in [9], with structural adoptions for FOA-domain. We adjust conditioning token dimensions to $512$ and introduce a projecting conditioning layer to meet dimensionality requirements. The model is trained using the v-objective approach [24] to minimize the Mean Squared Error (MSE) between the true noise sampled from the distribution $\epsilon\sim\mathcal{N}(0,\mathbf{I})$ and predicted noise $\epsilon_{\theta}(z_{t},t,c)$ at step $t$ with conditioning $c$, given the latent $z_{t}$ as follows:
$\mathcal{L}_{\text{Diffusion}}=\left\|\epsilon_{\theta}(z_{t},t,c)-\epsilon\right\|_{2}^{2}$ (2)
@@ 2.7 Evaluation Metrics
To quantify the difference between the ground truth and estimated directions of arrivals (DoA), we also include their spatial angle ${\Delta}_{\text{Angular}}$ [28] calculated using equations 6 and  7 as follows:
@@ 3 Results and Discussion
Table 3 presents the quality and spatial accuracy evaluations for both descriptive and parametric ImmerseDiffusion models. As shown in this table, the descriptive model significantly outperforms the parametric model in FADELSA score. However, this advantage may be influenced by the use of ELSA embeddings as a conditioner in the descriptive model, which the parametric model lacks. In terms of KL divergence and CLAP scores, on the other hand, the parametric model performs similarly to the descriptive model, slightly outperforming in one and under-performing in the other. This comparable performance underlines the effectiveness of the parametric training. In fact, even though the parametric model is trained on non-spatial prompts, it is capable of producing spatially plausible audio content thanks to the additional spatial parametrization.
@@
It is worth mentioning that the signal processing methods used for spatial adherence evaluation offer two key advantages: simplicity, which is a crucial criterion for evaluation metrics, and accessibility and generality, as they do not require pre-trained models or pose the risk of training data biases, unlike deep learning methods. However, due to higher lower-bound errors in some metrics (e.g., distance or spatial angle differences) as shown in Table 3, future extension of this work can include leveraging deep learning-based methods such as  [29] to improve measurement accuracy.
@@ 4 Conclusion
This paper introduced ImmerseDiffusion, an end-to-end model for generating 3D immersive soundscapes based on spatial, temporal, and environmental conditions. Using a spatial audio codec and a latent diffusion model, ImmerseDiffusion generates first-order ambisonics (FOA) audio. We presented two models: a descriptive model that creates spatial audio from text prompts detailing sound sources and contexts, and a parametric model that uses text prompts with explicit spatial and environmental parameters. Evaluations confirm that ImmerseDiffusion produces spatially accurate audio according to user-defined conditions.

### v2

@@ 1 Introduction
This paper introduces ImmerseDiffusion, a text-conditioned generative spatial audio model designed to produce 3D Ambisonic soundscapes with localization in horizontal, vertical, and distance dimensions, while also accounting for temporal and environmental factors like room size and reverberation. Our model generates FOA-domain audio, and features two operational modes: descriptive and parametric. The descriptive model generates spatial audio based on textual descriptions of sound sources and their spatial and environmental details, suited for narrative-driven applications such as cinematic audio. The parametric model combines textual descriptions of sound sources with numerical spatial parameters, ideal for machine-centered uses like game engines and virtual simulations. To the best of our knowledge, this is the first spatial audio generative model. Consequently, we introduce new metrics to evaluate the quality of the generated FOA audio. These metrics include Ambisonics Fréchet Audio Distance (FAD), spatial Kullback–Leibler (KL) divergence, and spatial CLAP scores. Additionally, we assess the spatial accuracy using azimuth, elevation, and distance L1 scores, besides spatial angle which are all based on sound intensity vectors [10].
@@ 2.2 Spatial Codec
To address compression, various approaches have been explored in the literature. Some models leverage autoencoders with quantized latents such as Vector Quantized Variational Autoencoders VQ-VAE (e.g., [12]) or Residual Vector Quantization RVQ-VAE (e.g., [13]). Other models leverage continuous VAE latents to avoid quantization information loss [2, 3, 9]. Following [3], we utilize the 1D convolutional U-Net autoencoder from the Descript Audio Codec (DAC) [14] and replace the discrete RVQ bottleneck with a continuous VAE bottleneck. We also eliminate the decoder’s $tanh()$ activation to prevent harmonic distortion, as noted in [9]. Additionally, we modify the training loss functions to suit the FOA components. Unlike stereo codecs, which leverage losses like Mid-Side Short Time Fourier Transform (MS-STFT) for proper stereo reconstruction, FOA does not need such losses due to the inherent orthogonality of the $X$, $Y$, and $Z$ channels. The training loss function of our FOA codec is as follows: {dmath} L_codec = λmrstft4 L_mrstft_W + λmrstft4 L_mrstft_X + λmrstft4 L_mrstft_Y + λmrstft4 L_mrstft_Z + λ_kl L_kl + λ_adv L_adv + λ_fm L_fm
In this context, $\mathcal{L}_{\text{mrstft}}$ represents the Multi-Resolution STFT loss for each channel, collectively contributing to the generation loss. $\mathcal{L}_{\text{kl}}$ denotes KL divergence loss applied to the VAE bottleneck, $\mathcal{L}_{\text{adv}}$ signifies the adversarial (discrimination) loss, while $\mathcal{L}_{\text{fm}}$ refers to the feature matching loss. The $\lambda$ parameters are weights perceptually assigned to adjust each corresponding loss contribution in total loss [3]. Our ambisonic codec model transforms a 4-channel FOA signal with length $L$ into $64$ channel latent with length $L/2048$ with an overall compression factor of $128$.
Table 1: Specifications and Quality and Spatial Accuracy Evaluation Results of the trained FOA Codecs on Spatial AudioCaps.
FOA Codec Model
@@
[…]
1D-Conv U-Net(32X) 1024 128 1.69 2.83 1.55 0.29 2.05 1.56
1D-Conv U-Net (64X) 1024 64 1.70 2.62 1.57 0.24 1.82 1.57
1D-Conv U-Net (128X) 2048 64 1.73 2.55 1.57 0.16 1.49 1.57
1D-Conv U-Net-$\beta$ (128X) 2048 64 2.33 4.13 1.43 0.15 1.53 1.51
Measurement Error 0.17 0.14 1.36 1.03
$\downarrow$
$\downarrow$
$\downarrow$
$\downarrow$
@@ 2.3 Spatial, Temporal and Environmental Conditioning
ImmerseDiffusion offers spatial control in audio generation via two models: descriptive and parametric. The descriptive model utilizes text-based prompts to define the sound source, its spatial attributes, and environmental acoustics. In contrast, the parametric model uses non-spatial text prompts only for sound source descriptions while providing numerical values for spatial and environmental parameters. Moreover, in line with previous stereo diffusion models  [3, 9], both models employ temporal conditioning, including start time and total time as well. In Figure 1, (I) and (II) illustrate how descriptive and parametric conditioning incorporate location and environmental parameters through text or a combination of text and numerical values, along with temporal conditioning.
@@
For the parametric model, we replace ELSA text encoder with the LAION CLAP [5] text encoder, which is solely trained on non-spatial prompts to validate parametric conditioning and avoid potential spatial bias from ELSA embeddings. Alongside text embeddings, this model includes three numerical spatial conditioning parameters (azimuth, elevation, distance) and two environmental acoustic conditioning parameters (room size, 30dB decay time). It also comprises two temporal conditioning parameters, namely, start time and total time, similar to the descriptive model.
@@ 2.4 Diffusion Model
Several generative audio latent diffusion models utilize 1D convolutional U-Net structures for their diffusion blocks (e.g., [2, 3]). However, the Diffusion Transformer (DiT), initially proposed for image generation [22], has recently been adapted for music generation [9, 23]. In our work, we employ a DiT that features a cascade of blocks that include self-attention, cross-attention, and gated MLP components with layer normalization and skip connections, similar to that used in [9], with structural adoptions for FOA-domain. We adjust conditioning token dimensions to $512$ and introduce a projecting conditioning layer to meet dimensionality requirements. The model is trained using the v-objective approach [24] to minimize the Mean Squared Error (MSE) between the true $v$ and the predicted $v_{\theta}(z_{t},t,c)$ at timestep $t$, conditioned on $c$ and given the latent variable $z_{t}$ as follows:
$\mathcal{L}_{\text{Diffusion}}=\left\|v_{\theta}(z_{t},t,c)-v\right\|_{2}^{2}$ (2)
@@ 2.7 Evaluation Metrics
To quantify the difference between the ground truth and estimated direction of arrival (DoA), we also include their spatial angle ${\Delta}_{\text{Angular}}$ [28] calculated using equations (6) and  (7) as follows:
@@ 3 Results and Discussions
Table 3 shows the quality and spatial accuracy evaluations for both descriptive and parametric ImmerseDiffusion models. The descriptive model significantly outperforms the parametric model in FADELSA score. However, this advantage may be influenced by the use of ELSA embeddings as a conditioner in the descriptive model, which the parametric model lacks. In terms of KL divergence and CLAP scores, on the other hand, the parametric model performs similarly to the descriptive model, slightly outperforming in one and under-performing in the other. This comparable performance underlines the effectiveness of the parametric training. In fact, even though the parametric model is trained on non-spatial prompts, it is capable of producing spatially plausible audio content thanks to the additional spatial parametrization.
@@
It is worth mentioning that the signal processing methods used for spatial adherence evaluation offer two key advantages: simplicity, which is a crucial criterion for evaluation metrics, and generality, as they do not require pre-trained models or pose the risk of training data biases, unlike deep learning methods. However, due to higher lower-bound errors in some metrics (e.g., distance or spatial angle differences) as shown in Table 3, future extension of this work can include leveraging deep learning-based methods such as  [29] to improve measurement accuracy.
@@ 4 Conclusion
This paper introduced ImmerseDiffusion, an end-to-end model for generating 3D immersive soundscapes based on spatial, temporal, and environmental conditions. Using a spatial audio codec and a latent diffusion model, ImmerseDiffusion generates first-order ambisonics (FOA) audio. We presented two models: a descriptive model that creates spatial audio from text prompts detailing sound sources and contexts, and a parametric model that uses text prompts with explicit spatial and environmental parameters. Experimental evaluations demonstrated that ImmerseDiffusion produces spatially accurate audio according to user-defined conditions.